OpenStudy (anonymous):

lim h -> 0 = ((1+h)^3-1)/h

OpenStudy (anonymous):

\[\lim h -> 0 = ((1+h)^3-1)/h\]

OpenStudy (anonymous):

is the problem

OpenStudy (anonymous):

Please show steps

OpenStudy (anonymous):

l'hospital

OpenStudy (anonymous):

\[\lim h-> 0 = ((1+h)(1+h)(1+h)-1)/h\] \[\lim h -> 0 = ((h ^{2}+2h+1)(1+h)-1)/h \] \[\lim h -> 0 = ((h ^{3}+3h ^{2}+3h+1)-1)/h\] \[\lim h -> 0 = (h ^{3}+3h ^{2}+3h)/h\] \[\lim -> 0 = h(h ^{2}+3h+3)/h\] \[\lim h -> 0 = h ^{2}+3h+3\] \[= 0^{2} +3(0) +3\] \[= 3\]